# Oriented matroids and type $\mathbb{A}$ cluster categories

### v2

@@ 4. Showing the bijection
We now show that stackable triangulations of these oriented matroids are in bijection with equivalence classes of maximal green sequences
@@ 4.1. Preliminary results
We first show the following key result, which shows the relation between our oriented matroid $\mathcal{M}_{T}$ and the choice of extriangulated structure on $\mathscr{C}_{n}$ coming from the cluster tilting object $\mathsf{T}$ corresponding to $T+1$.
@@
Note that by Section 2.2, we have that $\operatorname{Ext}_{\mathscr{C}_{n}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ due to the fact that $a\prec b\prec c\prec d$. By definition of the extriangulated structure defined by $\mathsf{T}$, we have that $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ if and only if the non-zero map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ factors through $\Sigma\mathsf{T}$. This is then the case precisely if there is a non-zero summand $\mathsf{O}_{X}$ of $\Sigma\mathsf{T}$ such that there is a non-zero composition $\mathsf{O}_{B}\to\mathsf{O}_{X}\to\Sigma\mathsf{O}_{A}$, since the map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ is unique up to scalar, by Section 2.2. By Lemma 4.2 and the description of $\Sigma$ from Section 2.2, this happens if and only if we have $b\preccurlyeq x\prec c\prec d\preccurlyeq y\prec a$, where $X=\{x,y\}$, and this is the case if and only if $\chi_{T}(a,b,c,d)=+1$. ∎
@@
Proof.
We argue using the bounded derived category $\mathscr{D}_{n}:=\mathscr{D}^{b}(K\mathbb{A}_{n})$. By [16, Lemma 6.6, Lemma 6.7, Proof of Proposition 6.1], the indecomposable objects of $\mathscr{D}_{n}$ may be labelled $\mathsf{U}_{i_{0}i_{1}}$ for $(i_{0},i_{1})\in\mathbb{Z}^{2}$ with $i_{0}+2\leqslant i_{1}$ and $i_{1}\leqslant i_{0}+n+1$, with $\operatorname{Hom}_{\mathscr{D}_{n}}(\mathsf{U}_{i_{0}i_{1}},\mathsf{U}_{j_{0}j_{1}})\neq 0$ if and only if $i_{0}-1<j_{0}<i_{1}-1<j_{1}<i_{0}+n+2$.
Suppose that $\mathsf{O}_{P}\to\mathsf{O}_{Q}$ and $\mathsf{O}_{Q}\to\mathsf{O}_{R}$ in $\mathscr{C}_{n}$ compose to give a non-zero morphism $\mathsf{O}_{P}\to\mathsf{O}_{R}$. Then, since $\mathscr{C}_{n}$ is an orbit category of $\mathscr{D}_{n}$, we can lift these morphisms to morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ which compose to give a non-zero map, where $\psi(\mathsf{U}_{P^{\prime}})=\mathsf{O}_{P}$, $\psi(\mathsf{U}_{Q^{\prime}})=\mathsf{O}_{Q}$, and $\psi(\mathsf{U}_{R^{\prime}})=\mathsf{O}_{R}$, with $\psi\colon\mathscr{D}_{n}\to\mathscr{C}_{n}$ the functor sending an object to its orbit. We have non-zero morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ composing to give a non-zero morphism $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{R^{\prime}}$ if and only if there are non-zero morphisms between these three pairs of objects, which is the case if and only if $p^{\prime}_{0}-1<q^{\prime}_{0}<p^{\prime}_{1}-1<q^{\prime}_{1}<p^{\prime}_{0}+n+2$, and $q^{\prime}_{0}-1<r^{\prime}_{0}<q^{\prime}_{1}-1<r^{\prime}_{1}<q^{\prime}_{0}+n+2$, and $p^{\prime}_{0}-1<r^{\prime}_{0}<p^{\prime}_{1}-1<r^{\prime}_{1}<p^{\prime}_{0}+n+2$. These three inequalities then all hold if and only if
(4.1) $p^{\prime}_{0}-1\leqslant q^{\prime}_{0}-1<r^{\prime}_{0}<p^{\prime}_{1}-1\leqslant q^{\prime}_{1}-1<r^{\prime}_{1}<p^{\prime}_{0}+n+2.$
By [16, Lemma 6.7], we have that $\psi(\mathsf{U}_{i^{\prime}_{0}i^{\prime}_{1}})=\mathsf{O}_{i_{0},i_{1}}$ where $\{i_{0},i_{1}\}=\{i^{\prime}_{0}\mod n+3,i^{\prime}_{1}\mod n+3\}$, and so (4.1) gives us the desired cyclic ordering $p_{0}-1\preccurlyeq q_{0}-1\prec r_{0}\prec p_{1}-1\preccurlyeq q_{1}-1\prec r_{1}$.
Conversely, if such a cyclic ordering holds, one can lift $\mathsf{O}_{P}$, $\mathsf{O}_{Q}$, and $\mathsf{O}_{R}$ to objects $\mathsf{U}_{P^{\prime}}$, $\mathsf{U}_{Q^{\prime}}$, and $\mathsf{U}_{R^{\prime}}$ of $\mathscr{D}_{n}$ such that (4.1) holds. This gives that the morphisms in $\mathscr{D}_{n}$ compose to give a non-zero morphism, and so the morphisms in $\mathscr{C}_{n}$ likewise compose to give a non-zero morphism. ∎
@@ 4.2. Maximal green sequence to triangulation
Proof.
We verify the properties of Definition 2.5. For (1), suppose that $\sigma,\tau\in\Delta_{\mathcal{G}}$. We have that $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$ are both the uniform rank-four oriented matroid on a 4-element set. Hence, all subsets of size three of $\sigma$ and $\tau$ are facets of these respective matroids and all subsets of $\sigma$ and $\tau$ are faces of $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$ respectively. It follows that $\sigma\cap\tau$ must be a common face of $\mathcal{M}_{\mathsf{T}}(\sigma)$ and $\mathcal{M}_{\mathsf{T}}(\tau)$.
@@
We now verify that (2) holds. Let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. Suppose that $\sigma\in\Delta_{\mathcal{G}}$. Indeed, let $\sigma=A_{i}\cup B_{i}$. Let $F$ be a facet of $\sigma$ containing $A_{i}$; one can argue similarly for facets of $\sigma$ containing $B_{i}$. If $F$ is a triangle of $T+1$, then it is a facet of $\mathcal{M}_{T}$ by Proposition 3.11.
@@
Thus, suppose that $F$ is not a triangle of $T+1$. Then, if $T_{j}$ is the triangulation corresponding to $\mathsf{T}_{j}$, let $T_{k}$ be the first triangulation where $F$ appears. By assumption, $k\neq 0$. There is then an exchange pair $(\mathsf{O}_{A_{k}},\mathsf{O}_{B_{k}})$ and a simplex $\sigma^{\prime}=A_{k}\cup B_{k}\in\Delta_{\mathcal{G}}$. Since $F$ is a triangle of $T_{k}$ but not of $T_{k-1}$, we have that $F$ must also be a facet of $\sigma^{\prime}$, and must contain $B_{k}$. This means that $\sigma\neq\sigma^{\prime}$, and so $F$ is contained in at least two simplices of $\Delta_{\mathcal{G}}$.
@@
We must now show that $F$ cannot be contained in more than two simplices of $\Delta_{\mathcal{G}}$. Suppose that we have another simplex $\sigma^{\prime\prime}\in\Delta_{\mathcal{G}}$ with $F$ as a facet of $\sigma^{\prime\prime}$. We have a total order on $\Delta_{\mathcal{G}}$ given by the order of the exchange pairs of $\mathcal{G}$. When a simplex $\tau$ corresponds to the exchange pair from $\mathsf{T}_{h-1}$ to $\mathsf{T}_{h}$, the lower facets of $\tau$ leave $T_{h-1}$ and are replaced by the upper facets of $\tau$. Since we have at least three simplices in $\Delta_{\mathcal{G}}$ which contain $F$, we have that $F$ enters the sequence of triangulations $T_{0},T_{1},\dots,T_{l}$, leaves the sequence, and then enters again. In particular, this must be true for some $1$-dimensional face $A$ of $F$, which we can choose to be an arc in the $(n+3)$-gon, since $n+3\geqslant 4$. This then gives us an indecomposable object $\mathsf{O}_{A}$ which enters $\mathcal{G}$, leaves, and then enters again. But this contradicts [11, Lemma 3.4].
@@
Finally, we must show (3): that no two simplices $\sigma,\tau\in\Delta$ overlap on a circuit. Suppose for contradiction that, in fact, there is a circuit $(C^{+},C^{-})$ of $\mathcal{M}_{T}$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq\underline{C}\setminus\{x\}$ for some $x\in C^{+}$. In particular, $\tau\supseteq C^{-}$. By rotating, we can assume that $\underline{C}=\{a\prec b\prec c\prec d\prec e\}$ with $C^{\pm}=\{a,c,e\}$ and $C^{\mp}=\{b,d\}$ by Proposition 3.9. Without loss of generality, suppose that $C^{+}=\{a,c,e\}$ and $C^{-}=\{b,d\}$. Suppose now that $\sigma$ precedes $\tau$ in the order given by $\mathcal{G}$. We have that $\chi_{\mathsf{T}}(b,c,d,e)=+1$ since the description of $C$ gives us that $\mathcal{M}_{T}(\underline{C})$ is the oriented matroid of a three-dimensional cyclic polytope with five vertices. Hence $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{ce},\mathsf{O}_{bd})\neq 0$ by Lemma 4.1. However, since $\sigma$ precedes $\tau$, we have that $\mathsf{O}_{ce}$ must precede $\mathsf{O}_{bd}$ in $\mathcal{G}$, which contradicts [11, Lemma 3.1(2)]. If $\tau$ precedes $\sigma$, then note that $\chi_{T}(a,b,c,d)=+1$, so $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{bd},\mathsf{O}_{ac})\neq 0$, which contradicts [11, Lemma 3.1(2)] in a similar way. ∎
@@
First note that $\mathcal{G}$ gives a total order $\leqslant_{\mathcal{G}}$ on its exchange pairs according to the order they appear. Since the exchange pairs of $\mathcal{G}$ are in bijection with the simplices of $\Delta_{\mathcal{G}}$, we have that $\leqslant_{\mathcal{G}}$ is also a total order on the simplices of $\Delta$. We claim that the relation $\rightsquigarrow$ on $\Delta$ is contained in the relation $\leqslant_{\mathcal{G}}$. From this it follows that $\rightsquigarrow$ is a partial order, since any reflexive transitive relation contained in a total order is a partial order.
@@
By Definition 3.13 and Lemma 4.1, we have that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ and $(\mathsf{O}_{A_{j}},\mathsf{O}_{B_{j}})$ are the respective exchange pairs corresponding to $\sigma$ and $\tau$. Let $\mathsf{T}_{j-1}$ be the cluster-tilting object of $\mathcal{G}$ on which the mutation of $\mathsf{O}_{A_{j}}$ for $\mathsf{O}_{B_{j}}$ is performed. Then, if $T_{j-1}$ is the triangulation of the $(n+3)$-gon corresponding to $\mathsf{T}_{j-1}$, we have that $T_{j-1}$ contains the lower facets of $\tau$. Since one of the lower facets of $\tau$ is an upper facet of $\sigma$, and each upper facet of $\sigma$ contains $B_{i}$, we have that $T_{j-1}$ contains the arc $B_{i}$. Then [11, Lemma 3.4] ensures that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ occurs before $\mathsf{T}_{j-1}$, and so $\sigma\leqslant_{\mathcal{G}}\tau$. ∎

### v3

@@ 4. Showing the bijection
We now show that stackable triangulations of these oriented matroids are in bijection with equivalence classes of maximal green sequences. Throughout this section we fix an integer $n\geqslant 1$.
@@ 4.1. Preliminary results
We first show the following key result, which shows the relation between our oriented matroid $\mathcal{M}_{T}$ and the choice of extriangulated structure on $\mathscr{C}_{n}$ coming from the cluster tilting object $\mathsf{T}$ corresponding to $T+1$. This lemma is the motivation for Definition 3.1.
@@
Note that by Section 2.2, we have that $\operatorname{Ext}_{\mathscr{C}_{n}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ due to the fact that $a\prec b\prec c\prec d$, since, by definition, $\operatorname{Ext}_{\mathscr{C}_{n}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})=\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\Sigma\mathsf{O}_{A})$. In the extriangulated structure defined by $\mathsf{T}$, we have that $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{B},\mathsf{O}_{A})\neq 0$ if and only if the non-zero map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ factors through $\Sigma\mathsf{T}$. This is then the case precisely if there is a non-zero summand $\mathsf{O}_{X}$ of $\Sigma\mathsf{T}$ such that there is a non-zero composition $\mathsf{O}_{B}\to\mathsf{O}_{X}\to\Sigma\mathsf{O}_{A}$, since the map $\mathsf{O}_{B}\to\Sigma\mathsf{O}_{A}$ is unique up to scalar, by Section 2.2. By Lemma 4.2 and the description of $\Sigma$ from Section 2.2, this happens if and only if we have $b\preccurlyeq x\prec c\prec d\preccurlyeq y\prec a$, where $X=\{x,y\}$, and this is the case if and only if $\chi_{T}(a,b,c,d)=+1$. ∎
Here we used the following lemma, whose proof we will postpone to the appendix.
@@ 4.2. Maximal green sequence to triangulation
First, we note that this triangulation respects equivalence of maximal green sequences.
@@
Let $T_{0},T_{1},\dots,T_{l}$ be the sequence of triangulations of the $(n+3)$-gon corresponding to the maximal green sequence $\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l}$. Let $\sigma=A_{i}\cup B_{i}$ be the simplex corresponding to the exchange pair $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ from $\mathsf{T}_{i-1}$ to $\mathsf{T}_{i}$. Then the triangulation $T_{i-1}$ is obtained from the triangulation $T_{i}$ by replacing the arc $A_{i}$ with the arc $B_{i}$. Hence $T_{i-1}$ and $T_{i}$ contain the quadrilateral $A_{i}\cup B_{i}$ and one can equivalently say that $T_{i}$ is obtained from $T_{i-1}$ by replacing the lower facets of $\sigma$ by the upper facets of $\sigma$. In this case, we say that $T_{i}$ is the increasing flip of $T_{i-1}$ given by $\sigma$. This perspective will be useful in the proof of the following proposition.
@@
In order to prove this proposition, we will rely on the following lemmas on maximal green sequences from [GW23].
Lemma 4.7 ([GW23, Lemma 3.4]).
Let $\mathsf{T}\in\mathscr{C}_{n}$ be a basic cluster-tilting object and let $\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l}$ be a maximal green sequence $\mathcal{G}$ of $\mathsf{T}$.
Then, given an indecomposable object $\mathsf{O}\in\mathscr{C}_{n}$, there exist $j,k\in[l]$ with $j\leqslant k$ such that
$\{\mathsf{T}_{i}:\mathsf{O}\text{ is a direct summand of }\mathsf{T}_{i}\}=\{\mathsf{T}_{j},\mathsf{T}_{j+1},\dots,\mathsf{T}_{k}\},$
if this set is non-empty.
In other words, an indecomposable object $\mathsf{O}$ enters $\mathcal{G}$ at most once, leaves $\mathcal{G}$ at most once, and enters before it leaves if it does both.
Lemma 4.8 ([GW23, Corollary 3.2(2)]).
Let $\mathsf{T}\in\mathscr{C}_{n}$ be a basic cluster-tilting object and let $\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l}$ be a maximal green sequence $\mathcal{G}$ of $\mathsf{T}$.
Let $\mathsf{X},\mathsf{Y}\in\mathscr{C}_{n}$ be indecomposable objects and suppose that there exist $j,k\in[l]$ with $j\leqslant k$ such that $\mathsf{X}$ is a direct summand of $\mathsf{T}_{j}$ and $\mathsf{Y}$ is a direct summand of $\mathsf{T}_{k}$. In this case, we say that $\mathsf{X}$ precedes $\mathsf{Y}$ in $\mathcal{G}$, and we have that $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{X},\mathsf{Y})=0$.
Proof of Proposition 4.6.
@@
We verify the properties of Definition 2.6. We first verify that (1) holds. Let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}=\mathsf{T}_{l}$, so that $T+1$ is the triangulation corresponding to $\mathsf{T}=\mathsf{T}_{0}$. Suppose that $\sigma\in\Delta_{\mathcal{G}}$, so that $\sigma=A_{i}\cup B_{i}$ for $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})\in\mathrm{Exch}(\mathcal{G})$. Let $F$ be a facet of $\sigma$ containing $A_{i}$; one can argue similarly for facets of $\sigma$ containing $B_{i}$. If $F$ is a triangle of $T+1$, then it is a facet of $\mathcal{M}_{T}$ by Proposition 3.12.
@@
Thus, suppose that $F$ is not a triangle of $T+1$. Then, if $T_{j}$ is the triangulation corresponding to $\mathsf{T}_{j}$, let $T_{k}$ be the first triangulation where $F$ appears. By assumption, $k\neq 0$. There is then an exchange pair $(\mathsf{O}_{A_{k}},\mathsf{O}_{B_{k}})$ and a simplex $\sigma^{\prime}=A_{k}\cup B_{k}\in\Delta_{\mathcal{G}}$. Since $F$ is a triangle of $T_{k}$ but not of $T_{k-1}$, we have that $F$ must also be a facet of $\sigma^{\prime}$, and must contain $B_{k}$. This means that $\sigma\neq\sigma^{\prime}$, since we cannot have $(A_{i},B_{i})=(A_{k},B_{k})$, as $F$ contains $A_{i}$ and $B_{k}$. Hence, $F$ is contained in at least two simplices of $\Delta_{\mathcal{G}}$.
@@
We must now show that $F$ cannot be contained in more than two simplices of $\Delta_{\mathcal{G}}$. Suppose that we have another simplex $\sigma^{\prime\prime}\in\Delta_{\mathcal{G}}$ with $F$ as a facet of $\sigma^{\prime\prime}$. We have a total order on $\Delta_{\mathcal{G}}$ given by the order of the exchange pairs of $\mathcal{G}$. When a simplex $\tau$ corresponds to the exchange pair from $\mathsf{T}_{h-1}$ to $\mathsf{T}_{h}$, the lower facets of $\tau$ leave $T_{h-1}$ and are replaced by the upper facets of $\tau$. Since we have at least three simplices in $\Delta_{\mathcal{G}}$ which contain $F$, we have that $F$ enters the sequence of triangulations $T_{0},T_{1},\dots,T_{l}$, leaves the sequence, and then enters again. In particular, this must be true for some $1$-dimensional face $A$ of $F$, which we can choose to be an arc in the $(n+3)$-gon, since $n+3\geqslant 4$. This then gives us an indecomposable object $\mathsf{O}_{A}$ which enters $\mathcal{G}$, leaves, and then enters again. But this contradicts Lemma 4.7.
@@
Finally, we must show (2): that no two simplices $\sigma,\tau\in\Delta$ overlap on a circuit. Suppose for contradiction that, in fact, there is a circuit $(C^{+},C^{-})$ of $\mathcal{M}_{T}$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq\underline{C}\setminus\{x\}$ for some $x\in C^{+}$. In particular, $\tau\supseteq C^{-}$. By rotating, we can assume that $\underline{C}=\{a\prec b\prec c\prec d\prec e\}$ with $C^{\pm}=\{a,c,e\}$ and $C^{\mp}=\{b,d\}$ by Proposition 3.9. Without loss of generality, suppose that $C^{+}=\{a,c,e\}$ and $C^{-}=\{b,d\}$. Suppose now that $\sigma$ precedes $\tau$ in the order given by $\mathcal{G}$. We have that $\chi_{\mathsf{T}}(b,c,d,e)=+1$ since the description of $C$ gives us that $\mathcal{M}_{T}(\underline{C})$ is the oriented matroid of a three-dimensional cyclic polytope with five vertices. Hence $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{ce},\mathsf{O}_{bd})\neq 0$ by Lemma 4.1. However, since $\sigma$ precedes $\tau$, we have that $\mathsf{O}_{ce}$ must precede $\mathsf{O}_{bd}$ in $\mathcal{G}$, which contradicts Lemma 4.8. If $\tau$ precedes $\sigma$, then note that $\chi_{T}(a,b,c,d)=+1$, so $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{O}_{bd},\mathsf{O}_{ac})\neq 0$, which contradicts Lemma 4.8 in a similar way. ∎
@@
First note that $\mathcal{G}$ gives a total order $\leqslant_{\mathcal{G}}$ on its exchange pairs according to the order in which they appear. Since the exchange pairs of $\mathcal{G}$ are in bijection with the simplices of $\Delta_{\mathcal{G}}$, we have that $\leqslant_{\mathcal{G}}$ is also a total order on the simplices of $\Delta$. We claim that the relation $\rightsquigarrow$ on $\Delta$ is contained in the relation $\leqslant_{\mathcal{G}}$. From this it follows that $\rightsquigarrow$ is a partial order, since any reflexive transitive relation contained in a total order is a partial order.
@@
By Definition 3.14 and Lemma 4.1, we have that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ and $(\mathsf{O}_{A_{j}},\mathsf{O}_{B_{j}})$ are the respective exchange pairs corresponding to $\sigma$ and $\tau$. Let $\mathsf{T}_{j-1}$ be the cluster-tilting object of $\mathcal{G}$ on which the mutation of $\mathsf{O}_{A_{j}}$ for $\mathsf{O}_{B_{j}}$ is performed. Then, if $T_{j-1}$ is the triangulation of the $(n+3)$-gon corresponding to $\mathsf{T}_{j-1}$, we have that $T_{j-1}$ contains the lower facets of $\tau$. Since one of the lower facets of $\tau$ is an upper facet of $\sigma$, and each upper facet of $\sigma$ contains $B_{i}$, we have that $T_{j-1}$ contains the arc $B_{i}$. This means that $\mathsf{O}_{B_{i}}$ is a direct summand of $\mathsf{T}_{j-1}$, and so Lemma 4.7 ensures that $(\mathsf{O}_{A_{i}},\mathsf{O}_{B_{i}})$ occurs before $\mathsf{T}_{j-1}$, which gives that $\sigma\leqslant_{\mathcal{G}}\tau$. ∎